OpenStudy (anonymous):

If (a+1/b)=4, find value of a3+1/b3

OpenStudy (agent0smith):

Again... this has infinite solutions, just like your question from yesterday... eg a=2, b=0.5 \[\large (2+1/0.5) = 4\]\[\large 2^3+1/0.5^3 = 16\] a=5, b=-1 \[\large (5+1/-1) = 4\]\[\large 5^3+1/-1^3 = 124\]